# TCLC-GS: Tightly Coupled LiDAR-Camera Gaussian Splatting for Surrounding Autonomous Driving Scenes

### v1

@@ Abstract
LiDAR-Camera Gaussian Splatting Real-time Rendering Surrounding Driving Views
@@ 1 Introduction
Urban-level reconstruction and rendering present significant challenges due to the vast scale of the unbounded environments and the sparse nature of the captured data. Fortunately, in autonomous vehicle settings, data from various modalities captured by multiple sensors are typically available. However, fully leveraging different modality data from multi-sensors for precise modelling and real-time rendering in urban scenes remains an open question in the field.
@@
In contrast to Neural Radiance Fields (NeRF), 3D Gaussian Splatting (3D-GS) [9], utilizes a more extensive 3D Gaussian representation for scenes, which not only achieves faster training but also facilitates real-time rendering. The initial 3D-GS method sought to initialize Gaussians using the points from structure-from-motion (SfM). However, this approach faces challenges in unbounded urban scenes within autonomous driving contexts, particularly when viewpoints are sparse. This sparsity can cause SfM techniques to fail in accurately and completely recovering scene geometries. To facilitate better 3D Gaussian initialization, pioneering research [27, 25, 5] has introduced LiDAR priors into the 3D-GS process, enabling more accurate geometries and ensuring rendering consistency of multiple surrounding views. However, directly initializing 3D Gaussians’ positions with LiDAR points does not thoroughly exploit the rich 3D geometric information embedded within 3D LiDAR points, such as depth and geometric features.
@@ 2 Related Work
NeRF-based techniques [14, 2, 22] have demonstrated significant effectiveness in large-scale urban scenarios for autonomous driving. NeRF-W [14] integrates frame-specific codes within its rendering pipeline, effectively managing photometric variations and transient objects. Mip-NeRF 360 [2], an advancement of Mip-NeRF [1], adapts to unbounded scenes by compressing the entire space into a bounded area, thereby enhancing the representativeness of position encoding. Block-NeRF [22] tackles the challenge of modeling large-scale outdoor scenes by partitioning the target urban environment into several blocks, with each segment represented by an individual NeRF network, significantly boosting the overall modeling capability.
@@ 3 Methodology
Overview: The pipeline for our TCLC-GS is detailed in Fig. 2. The TCLC-GS framework is composed of two primary learning components: 1) the octree implicit feature with SDF and RGB decoders (Fig. 2 upper part), and 2) 3D Gaussians with depth and RGB splattings (Fig. 2 lower part). The LiDAR and camera data are tightly integrated in an uniform framework.
@@
3.1 Problem Definition
In an urban street scene, given a sequence of surrounding images and LiDAR data collected from a vehicle-mounted system, our goal is to develop a model to reconstruct the environment and generate photorealistic images and depths from novel viewpoints. A set of surrounding images $\mathcal{I}$ are captured by multiple surrounding cameras with corresponding intrinsic matrices $I$ and extrinsic matrices $E$. A set of 3D points $\mathcal{P}$ captured by LiDAR with corresponding extrinsic matrices $E^{\prime}$. Meantime, the vehicle trajectory $\mathcal{T}$ is provided or calculated by multi-sensor-based odometry. Our model $\mathcal{F}$ aims to achieve precise 3D reconstruction and synthesize novel viewpoints at novel camera pose $[E_{n},I_{n}]$ by rendering ^$\hat{\mathcal{I}},\hat{\mathcal{D}}=\mathcal{F}(E_{n},I_{n})$.
3.2 Hierarchical Octree Implicit Feature
To encapsulate fine-grained geometric details and contextual structure information of the global 3D scene, we learn and store implicit features through learnable octree-based hierarchical feature grids $\mathcal{O}$ similar to NGLOD [21] from a sequence of LiDAR data. The implicit features can be further decoded into signed distance values (SDFs) and RGB colors through a shallow dual-branch MLP decoder.
We first merge a sequence of LiDAR sweeps using vehicle trajectory $\mathcal{T}$ to obtain the global point cloud of the scene. We construct a $L$ level octree using the point clouds, and store latent features at eight corners per octree node in each tree level. Multi-resolution features are stored across different octree levels, where the octree feature at level $i$ is denoted as $O_{i}\in\mathcal{O}$. The fine-grained features at the lower levels of the octree capture detailed geometry, whereas the coarse features at the higher levels encode broader contextual information. To optimize the implicit features, we sample 3D query points $p\in\mathbb{R}^{3}$ within the truncation region of beam endpoints along the LiDAR ray. For a query point $p$, we compute its feature vector $O_{i}(p)$ by trilinear interpolating its corresponding features at octree node corners. We feed both position encoding and multi-resolution octree features retrieved by a query point $p$ into the dual branch MLP decoder, i.e., SDF decoder $D_{s}$ to estimate the SDF $\hat{s}_{p}$, and RGB decoder $D_{c}$ to estimate the RGB color $\hat{c}_{p}$,
$\hat{s}_{p}=D_{s}(f(p),O_{i}(p)),\hat{c}_{p}=D_{c}(f(p),O_{i}(p)),i=1,2,...L$ (1)
@@ 3.2 Hierarchical Octree Implicit Feature
$\mathcal{L}_{eik}(p)=(1-||\nabla D_{s}(f(p),O_{i}(p))||)^{2},$ (3)
$\mathcal{L}_{smooth}(p)=||\nabla D_{s}(f(p),O_{i}(p))-\nabla D_{s}(f(p+\epsilon),O_{i}(p+\epsilon))||^{2},$ (4)
@@
After learning and storing the octree implicit features, which includes both geometry and RGB information, we can generate a colorized 3D mesh. Specifically, we first uniformly sample 3D points within the scene’s 3D spatial space. Each of these points is then used to retrieve octree implicit features, which are then fed into a dual-branch decoder. Subsequently, we generate a colorized 3D mesh $\mathcal{M}$ in the form of a triangle mesh by marching cubes [12] based on the decoder predicted SDFs and colors. Furthermore, from this 3D mesh $\mathcal{M}$, we generate a set of dense depth image $\mathcal{D}$ along the vehicle trajectory by Ray Tracing [6] as,
@@
Through representing the scene as a mesh, the learning of 3D Gaussians is enhanced with a more robust geometry and appearance prior. In detail, the mesh fully exploits the geometries from LiDAR points in terms of continuty/completion and density. Compared with sparse raw LiDAR points, the mesh representations are able to reconstruct missed geometries due to sparse scans and viewpoints, and provides scalable and adaptive geometry priors. Meanwhile, the colorized 3D mesh provide RGB information for any 3D point positions sampled on its surface. This is a significant improvement over the intuitive method of projecting LiDAR points onto surrounding images, which typically assigns colors to only about one-third of the LiDAR points due to limited sensor overlap between camera and LiDAR. In contrast to the sparse depth obtained from LiDAR points, the dense depth rendered from the 3D mesh provides more robust supervision and alleviate the over-fitting during training.
3.4 LiDAR-Camera Gaussian Splatting
3D-GS [9] employs a collection of 3D Gaussians to represent a scene. It uses a tile-based rasterization process, enabling real-time alpha blending of multiple Gaussians. Different from directly initializing Gaussian by LiDAR points, our method constructs a 3D Gaussians $\mathcal{GS}$ according to the vertex $v$ of the 3D mesh $\mathcal{M}$. Each vertex $v$ provides position $\mu$, color $c$, and the vertex-retrieved octree implicit feature $O(v)$ to initialize the 3D Gaussians’ attributes as,
$\mathcal{GS}=\{\mu(v),s(v),q(v),c(v)|\mathcal{M};o(O_{i}(v)),sh(c,O_{i}(v)),O_{i}(v)|\mathcal{M},\mathcal{O}\},i=1,..L.$ (7)
The 3D Gaussians’ attributes include mean position $\mu\in\mathbb{R}^{3}$, anisotropic covariance $\Sigma\in\mathbb{R}^{3\times 3}$ represented by a scale $s\in\mathbb{R}^{3}$ (diagonal matrix) and rotation $q$ (unit quaternion), opacity $o$ and spherical harmonics coefficients $sh$. The vertices and colors of 3d mesh $\mathcal{M}$ serve as geometry and appearance priors for $\mu,s,q,c$ and $o,sh$ properties of 3D Gaussians. Additionally, the octree implicit features are also incorporated to improve appearance descriptions $o,sh$ of 3D Gaussians. Please note that the newly added properties $O_{i}(v)$ are dynamically controlled through cloning, and pruning alongside other attributes by adaptive density control during optimization. Additionally, to enhance training robustness, we use very shallow MLPs to encode the $c$ and $O_{i}(v)$ before their integration into the 3D Gaussians.
@@ 3.4 LiDAR-Camera Gaussian Splatting
By sorting the Gaussians according to their depth within the camera space, we can effectively query the attributes of each 2D Gaussian. This step facilitates the following volume rendering process to estimate the color $C$ and depth $D$ of each pixel as,
$C=\sum_{i=1}^{N}T_{i}\alpha_{i}c_{i},D=\sum_{i=1}^{N}T_{i}\alpha_{i}z_{i},T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (9)
where $z$ denotes the distance from the image plane to the Gaussian point center. $\alpha$ is calculated by multiplying the opacity $o$ with the probability of the 2D Gaussian according to $\Sigma^{\prime}$.
Given a specific view ($E,I$) in a scene containing $N$ 3D Gaussians, both the image $\hat{\mathcal{I}}$ and depth $\hat{\mathcal{D}}$ are rendered by the differentiable rendering function $\mathcal{F}$ as,
^$\hat{\mathcal{I}},\hat{\mathcal{D}}=\mathcal{F}(\{\mathcal{GS}_{i},i=1,2,..N\}|E,I).$ (10)
The parameters of $\mathcal{GS}$ will be optimized during training by the objective function $\mathcal{L}$ as following,
$\mathcal{L}=(1-\lambda_{ssim})\mathcal{L}_{1}(\mathcal{I},\hat{\mathcal{I}})+\lambda_{ssim}\mathcal{L}_{ssim}(\mathcal{I},\hat{\mathcal{I}})+\lambda_{depth}\mathcal{L}_{1}(\mathcal{D},\hat{\mathcal{D}}),$ (11)
where $\lambda_{ssim}$ and $\lambda_{depth}$ are scale factors, and $\mathcal{L}_{ssim}$ refers to Structural Similarity Index Measure (SSIM) term. $\mathcal{I}$ refers to the original RGB image, and $\mathcal{D}$ refers to the dense depth rendered from 3D mesh by Ray Tracing.
In order to mitigate the scale ambiguity due to the long duration of the driving scenario, we employ the incremental optimization strategy [27] which involves incrementally adding 3D Gaussians by uniformly segmenting the scenario into a series of bins based on the LiDAR depth range. The enhancement of Gaussian Splatting optimization is achieved by improving the initialization of 3D Gaussians’ geometry and appearance, enriching the appearance descriptions of 3D Gaussians, and integrating dense depth supervision.
@@ 4.1 Datasets, Metrics and Baselines
4.2 Evaluation on Waymo Open Dataset
We evaluated the proposed method by comparing it with the baseline on the Waymo Open dataset. The performance comparison of image and depth synthesis from novel views, relative to the main baseline, is detailed in Table 1 and Table 2 separately. Additional comparisons of image synthesis from novel views against a broader range of baselines are depicted in Table 3. As indicated in Table 1, our method outperforms 3D-GS across individual scenes in terms of PSNR, SSIM, and LPIPS metrics for novel image synthesis. Furthermore, Table 3 shows that our method exceeds a broader array of baselines in average performance across these same metrics (PSNR, SSIM, and LPIPS) for novel image synthesis. Similarly, Table 2 demonstrates that our approach significantly surpasses 3D-GS in the depth synthesis metrics of AbsRel, RSME, and RSMElog. The significant improvement in depth synthesis performance can be attributed to the robust supervision provided by the rendered dense depths derived from the generated accurate 3D mesh. For a visual perspective, the comparison results of image synthesis from novel surrounding views are illustrated in Fig. 4. We can see that TCLC-GS renders more clear and accurate RGB images than the 3D-GS, especially on roadside objects and in distant areas viewed from the front, front-left, and front-right perspectives. Similarly, the visual comparison results for depth synthesis from novel surrounding views are showcased in Fig. 4. Here, TCLC-GS is observed to render denser and sharper depths compared to 3D-GS, especially in areas further away in the front, front-left, and front-right views.
@@ 4.2 Evaluation on Waymo Open Dataset
4.3 Evaluation on nuScenes Dataset
We conducted further comprehensive evaluations comparing our method with the baseline on the nuScenes dataset. The results of image and depth synthesis from novel viewpoints, compared with the primary baseline 3D-GS, are detailed in Table 4 and Table 5 separately. According to Table 4, our method outperforms 3D-GS in the metrics of PSNR, SSIM, and LPIPS for novel image synthesis. Likewise, as indicated in Table 5, our approach significantly excels beyond 3D-GS in the depth synthesis metrics of AbsRel, RSME, and RSMElog. The lower depth performance of our method observed on the nuScenes dataset, compared to the Waymo dataset, is due to the nuScenes dataset’s reliance on relatively sparse 32-line LiDAR data. In contrast, the Waymo dataset includes much denser LiDAR data, which contributes to a more accurate generation of 3D mesh and, subsequently, better rendered dense depth. Visual comparison results depicted in Figure 5 and 6 illustrate that our method not only produces clearer synthesized RGB images with fewer artifacts and less blurring but also generates sharper and denser synthesized depth images compared to the 3D-GS baseline.
Sequence 3D-GS [9] PSNR↑/SSIM↑/LPIPS↓ TCLC-GS PSNR↑/SSIM↑/LPIPS↓
Scene-0008 25.32/0.82/0.26 26.41/0.85/0.22
Scene-0051 25.84/0.85/0.25 26.78/0.86/0.23
Scene-0058 24.71/0.81/0.27 26.45/0.88/0.20
Scene-0062 25.91/0.87/0.22 27.47/0.89/0.17
Scene-0129 26.87/0.88/0.21 28.65/0.90/0.16
Scene-0382 23.96/0.82/0.32 25.76/0.84/0.26
Average 25.44/0.84/0.26 26.92/0.87/0.21
Scene-0008 0.31/9.97/0.47 0.07/4.27/0.12
Scene-0051 0.47/8.68/0.60 0.05/2.21/0.07
Scene-0058 0.35/10.48/0.55 0.06/4.23/0.11
Scene-0062 0.32/8.22/0.39 0.04/4.03/0.08
Scene-0129 0.29/8.01/0.44 0.05/2.83/0.09
Scene-0382 0.32/10.91/0.51 0.05/4.58/0.10
Average 0.34/9.38/0.49 0.05/3.69/0.10
4.4 Ablation Study
To demonstrate the individual effectiveness of each component in our method, we conducted ablation studies using the Waymo dataset. Since our contributions mainly include colorized 3D mesh, octree implicit representation, and dense depth supervision, our ablation study analyses the impact of our designs from these aspects. We train five different variations: 1) TCLC-GS without 3D mesh, initializing 3D Gaussian using original 3D LiDAR points; 2) TCLC-GS without colorized 3D mesh, initializing 3D Gaussian using 3D mesh without color information; 3) TCLC-GS without octree implicit representation, initializing 3D Gaussian using colorized 3D mesh without octree representation; 4) TCLC-GS without dense depth supervision, training TCLC-GS only using RGB supervision; 5) TCLC-GS full method. The average values of the evaluation metrics across six testing scenes from the Waymo dataset are given in Table 6. These ablation results indirectly validate the effectiveness of colorized 3D mesh, octree implicit representation, and dense depth supervision, underscoring their contributions to the overall performance.
@@ 4.4 Ablation Study
4.5 Running-time Analysis
We provide the running time performance evaluation of TCLC-GS tested on the Waymo and nuScenes datasets. Utilizing a single NVIDIA GeForce RTX 3090 Ti, TCLC-GS attains real-time rendering speeds with high-resolution images, achieving around 90 FPS with an image resolution of 1920$\times$1280 on the Waymo dataset, and around 120 FPS with an image resolution of 1600$\times$900 on the nuScenes dataset. The total training time for one scene in Waymo or nuScenes datasets takes approximately five hours on the same GPU, with the duration depending on the sequence length of the scene. Our solution enhances the quality of fast 3D modeling and real-time RGB/depth rendering, maintaining the inherent efficiency of Gaussian Splatting.
5 Conclusion
6 Supplementary materials
In this Appendix, we firstly provide more details of our TCLC-GS implementation, including network architecture, and hyper-parameters of octree feature grid and 3D Gaussians in Appendix 7. Secondly, we provide performance comparison with additional baselines on the nuScenes dataset in Appendix 8. Thirdly, more comparison visualizations with the baseline are presented in Appendix 9. Fourthly, the runtime analysis is given in in Appendix 10. Lastly, three video demos are attached in Appendix 11.
7 Implementation Details
8 Additional Baselines Evaluation
In Table 7, we present a performance comparison between TCLC-GS and additional baselines on the nuScenes dataset, showcasing the average performance on PSNR, SSIM, and LPIPS metrics for novel image synthesis. Our method surpasses these baselines in all mentioned metrics, demonstrating its superior performance in image synthesis.
Method PSNR↑ SSIM↑ LPIPS↓
Mip-NeRF [1] 18.10 0.63 0.46
Mip-NeRF 360 [2] 23.45 0.74 0.35
S-NeRF [10] 25.62 0.77 0.27
3D-GS [9] 25.44 0.84 0.26
TCLC-GS 26.92 0.87 0.21
9 More Visualization Results
Visualization Results on the Waymo Open dataset: We present visual comparison results of image and depth synthesis on the Waymo dataset in Fig. 8, Fig. 10 and Fig. 12. It can be seen that TCLC-GS produces clearer images compared to 3D-GS [9]. Additionally, the depth images generated by TCLC-GS are denser and sharper than those produced by 3D-GS.
10 Runtime Analysis
We present a running time performance comparison between our TCLC-GS and various baselines on the Waymo and nuScenes datasets, detailed in Table 8 and Table 9, respectively. Compared to NeRF-based methods such as SUDS [23] and S-NeRF [10], TCLC-GS significantly outperforms the speed of these methods, achieving real-time running performance. When compared to 3D-GS [9], TCLC-GS offers higher accuracy in both RGB and depth rendering, and still maintains the performance efficiency.

### v2

# TCLC-GS: Tightly Coupled LiDAR-Camera Gaussian Splatting for Autonomous Driving

@@ Abstract
LiDAR-Camera Gaussian Splatting Real-time Rendering Sensor Fusion Autonomous Driving
@@ 1 Introduction
Urban-level reconstruction and rendering present significant challenges due to the vast scale of the unbounded environments and the sparse nature of the captured data. Fortunately, in autonomous vehicle settings, data from various modalities captured by multiple sensors are typically available. However, fully leveraging different modality data from multi-sensors for precise modeling and real-time rendering in urban scenes remains an open question in the field.
@@
In contrast to Neural Radiance Fields (NeRF), 3D Gaussian Splatting (3D-GS) [11], utilizes a more extensive 3D Gaussian representation for scenes, which not only achieves faster training but also facilitates real-time rendering. The initial 3D-GS method sought to initialize Gaussians using the points from structure-from-motion (SfM). However, this approach faces challenges in unbounded urban scenes within autonomous driving contexts, particularly when viewpoints are sparse. This sparsity can cause SfM techniques to fail in accurately and completely recovering scene geometries. To facilitate better 3D Gaussian initialization, pioneering research [31, 28, 5] has introduced LiDAR priors into the 3D-GS process, enabling more accurate geometries and ensuring rendering consistency of multiple surrounding views. However, directly initializing 3D Gaussians’ positions with LiDAR points does not thoroughly exploit the rich 3D geometric information embedded within 3D points, such as depth and geometric features.
@@ 2 Related Work
NeRF-based techniques [16, 2, 24] have demonstrated significant effectiveness in large-scale urban scenarios for autonomous driving. NeRF-W [16] integrates frame-specific codes within its rendering pipeline, effectively managing photometric variations and transient objects. Mip-NeRF 360 [2], an advancement of Mip-NeRF [1], adapts to unbounded scenes by compressing the entire space into a bounded area, thereby enhancing the representativeness of position encoding. Block-NeRF [24] tackles the challenge of modeling large-scale outdoor scenes by partitioning the target urban scene into several blocks, with each segment represented by an individual NeRF network, boosting the overall modeling capability.
@@ 3 Methodology
Overview: The pipeline for our TCLC-GS is detailed in Fig. 2. The TCLC-GS framework is composed of two primary learning components: 1) the octree implicit feature with SDF and RGB decoders (Fig. 2 upper part), and 2) 3D Gaussians with depth and RGB splattings (Fig. 2 lower part). The LiDAR and camera data are tightly integrated in a uniform framework.
@@
Problem Definition: In an urban street scene, given a sequence of surrounding images and LiDAR data collected from a vehicle-mounted system, our goal is to develop a model to reconstruct the environment and generate photorealistic images and depths from novel viewpoints. A set of surrounding images $\mathcal{I}$ are captured by multiple surrounding cameras with corresponding intrinsic matrices $I$ and extrinsic matrices $E$. A set of 3D points $\mathcal{P}$ captured by LiDAR with corresponding extrinsic matrices $E^{\prime}$. Meantime, the vehicle trajectory $\mathcal{T}$ is provided or calculated by multi-sensor-based odometry. Our model $\mathcal{R}$ aims to achieve precise 3D reconstruction and synthesize novel viewpoints at novel camera pose $[E_{n},I_{n}]$ by rendering ^$\hat{\mathcal{I}},\hat{\mathcal{D}}=\mathcal{R}(E_{n},I_{n})$.
Hierarchical Octree Implicit Feature: To encapsulate fine-grained geometric details and contextual structure information of the global 3D scene, we learn and store implicit features through learnable octree-based hierarchical feature grids $\mathcal{F}$ similar to NGLOD [23] from a sequence of LiDAR data. The implicit features can be further decoded into signed distance values (SDFs) and RGB colors through a shallow dual-branch MLP decoder.
We first merge a sequence of LiDAR sweeps using vehicle trajectory $\mathcal{T}$ to obtain the global point cloud of the scene. We construct a $L$ level octree using the point clouds, and store latent features at eight corners per octree node in each tree level. Multi-resolution features are stored across different octree levels, where the octree feature at level $i$ is denoted as $F_{i}\in\mathcal{F}$. The fine-grained features at the lower levels of the octree capture detailed geometry, whereas the coarse features at the higher levels encode broader contextual information. To optimize the implicit features, we sample 3D query points $p\in\mathbb{R}^{3}$ within the truncation region of beam endpoints along the LiDAR ray. For a query point $p$, we compute its feature vector $F_{i}(p)$ by trilinear interpolating its corresponding features at octree node corners. We feed both position encoding and multi-resolution octree features retrieved by a query point $p$ into the dual branch MLP decoder, i.e., SDF decoder $D_{s}$ to estimate the SDF $\hat{s}_{p}$, and RGB decoder $D_{c}$ to estimate the RGB color $\hat{c}_{p}$,
$\hat{s}_{p}=D_{s}(f(p),F_{i}(p)),\hat{c}_{p}=D_{c}(f(p),F_{i}(p)),i=1,2,...L$ (1)
@@
$\mathcal{L}_{eik}(p)=(1-||\nabla D_{s}(f(p),F_{i}(p))||)^{2},$ (3)
$\mathcal{L}_{smooth}(p)=||\nabla D_{s}(f(p),F_{i}(p))-\nabla D_{s}(f(p+\epsilon),F_{i}(p+\epsilon))||^{2},$ (4)
@@
3D Mesh and Dense Depth Generation: After learning and storing the octree implicit features, which includes both geometry and RGB information, we can generate a colorized 3D mesh. Specifically, we first uniformly sample 3D points within the scene’s 3D spatial space. Each of these points is then used to retrieve octree implicit features, which are then fed into a dual-branch decoder. Subsequently, we generate a colorized 3D mesh $\mathcal{M}$ in the form of a triangle mesh by marching cubes [14] based on the decoder predicted SDFs and colors. Furthermore, from this 3D mesh $\mathcal{M}$, we generate a set of dense depth images $\mathcal{D}$ along the vehicle trajectory by Ray Tracing [6] as,
@@
Through representing the scene as a mesh, the learning of 3D Gaussians is enhanced with a more robust geometry and appearance prior. In detail, the mesh fully exploits the geometries from LiDAR points in terms of continuity/completion and density. Compared with sparse raw LiDAR points, the mesh representations are able to reconstruct missed geometries due to sparse scans and viewpoints, and provides scalable and adaptive geometry priors. Meanwhile, the colorized 3D mesh provides RGB information for any 3D point positions sampled on its surface. This is a significant improvement over the intuitive method of projecting LiDAR points onto surrounding images, which typically assigns colors to only about one-third of the LiDAR points due to limited sensor overlap between camera and LiDAR. In contrast to the sparse depth obtained from LiDAR points, the dense depth rendered from the 3D mesh provides more robust supervision and alleviate the over-fitting during training.
LiDAR-Camera Gaussian Splatting: 3D-GS [11] employs a collection of 3D Gaussians to represent a scene. It uses a tile-based rasterization process, enabling real-time alpha blending of multiple Gaussians. Different from directly initializing Gaussian by LiDAR points, our method initials 3D Gaussians $\mathcal{GS}$ on the 3D mesh $\mathcal{M}$. We follow the rule of Gaussians bound to the triangles [9] to keep the 3D Gaussians flat and aligned with the mesh triangles. Each vertex $v$ in the mesh provides position $\mu$, color $c$, and the vertex-retrieved octree implicit feature $F(v)$ to initialize the 3D Gaussians’ attributes as,
$\mathcal{GS}=\{\mu(v),s(v),q(v),c(v)|\mathcal{M};o(F_{i}(v)),sh(c(v),F_{i}(v)),F_{i}(v)|\mathcal{M},\mathcal{F}\},i=1,..L.$ (7)
The 3D Gaussians’ attributes include mean position $\mu\in\mathbb{R}^{3}$, anisotropic covariance $\Sigma\in\mathbb{R}^{3\times 3}$ represented by a scale $s\in\mathbb{R}^{3}$ (diagonal matrix) and rotation $q$ (unit quaternion), opacity $o$ and spherical harmonics coefficients $sh$. The vertices and colors of 3D mesh $\mathcal{M}$ serve as geometry and appearance priors for $\mu,s,q,c$ and $o,sh$ properties of 3D Gaussians. Additionally, the octree implicit features are also incorporated to improve appearance descriptions $o,sh$ of 3D Gaussians. Please note that the newly added properties $F_{i}(v)$ are dynamically controlled through cloning, and pruning alongside other attributes by adaptive density control during optimization. Additionally, to enhance training robustness, we use very shallow MLPs to encode the $c$ and $F_{i}(v)$ before their integration into the 3D Gaussians.
@@
By sorting the Gaussians according to their depth within the camera space, we can effectively query the attributes of each 2D Gaussian. This step facilitates the following volume rendering process to estimate the color $C$, depth $D$ and accumulated opacity $O$ of each pixel as,
$C=\sum_{i=1}^{N}T_{i}\alpha_{i}c_{i},D=\sum_{i=1}^{N}T_{i}\alpha_{i}z_{i},O=\sum_{i=1}^{N}T_{i}\alpha_{i},T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (9)
where $z$ denotes the distance from the image plane to the Gaussian point center. $\alpha$ is calculated by $\alpha_{i}=o_{i}\cdot exp(-\frac{1}{2}(x-\mu_{i})^{T}\Sigma^{{}^{\prime}-1}(x-\mu_{i}))$
Given a specific view ($E,I$) in a scene containing $N$ 3D Gaussians, both the image $\hat{\mathcal{I}}$ and depth $\hat{\mathcal{D}}$ are rendered by the differentiable rendering function $\mathcal{R}$ as,
^$\hat{\mathcal{I}},\hat{\mathcal{D}}=\mathcal{R}(\{\mathcal{GS}_{i},i=1,2,..N\}|E,I).$ (10)
Note all depths used in this paper are inverse depths. The parameters of $\mathcal{GS}$ will be optimized during training by the objective function $\mathcal{L}$ as following,
$\begin{split}\mathcal{L}=(1-\lambda_{ssim})\mathcal{L}_{1}(\mathcal{I},\hat{\mathcal{I}})+\lambda_{ssim}\mathcal{L}_{ssim}(\mathcal{I},\hat{\mathcal{I}})+\lambda_{depth}\mathcal{L}_{1}(\mathcal{D},\hat{\mathcal{D}})+\\ \lambda_{smooth}\mathcal{L}_{smooth}(\mathcal{I},\hat{\mathcal{D}})+\lambda_{sky}\mathcal{L}_{sky},\end{split}$ (11)
where $\lambda_{ssim}$, $\lambda_{depth}$, $\lambda_{smooth}$ and $\lambda_{sky}$ are scale factors. $\mathcal{I}$ refers to the original RGB image, and $\mathcal{D}$ refers to the dense depth rendered from 3D mesh by Ray Tracing. $\mathcal{L}_{ssim}$ refers to Structural Similarity Index Measure (SSIM) term between the original RGB and rendered RGB images, and $\mathcal{L}_{smooth}$ refers to inverse depth smooth [7] term between the original RGB image and rendered depth. $\mathcal{L}_{sky}$ refers to the sky opacity loss [29] defined as BCE loss,
$\mathcal{L}_{sky}=\mathcal{L}_{bce}(O,1-M_{sky})$ (12)
where $M_{sky}$ is the sky mask from a pretrained segmentation model SegFormer [26]. The $\mathcal{L}_{sky}$ drives the opacity of sky pixels toward zero to be transparent, which pushes the 3D Gaussian positions of the sky far away. The $\mathcal{L}_{sky}$ significantly reduces the sky artifacts, enabling the depth synthesis with clear sky areas. These sky masks are only necessary during training and not during inference.
The enhancement of Gaussian optimization is achieved by improving the initialization of 3D Gaussians’ geometry and appearance, enriching the appearance descriptions of 3D Gaussians, and integrating dense depth supervision. In order to mitigate the scale ambiguity due to the long duration of the driving scenario, we employ the incremental optimization strategy [31], which involves incrementally adding 3D Gaussians by uniformly segmenting the scenario into a series of bins based on the LiDAR depth range. For each bin, we further employ position-aware point adaptive control [5] during Gaussian optimization, which utilizes smaller points for nearby positions and larger points for distant locations in the unbounded scene.
@@ 4 Experiments
Sequence 3D-GS [11] PSNR↑/SSIM↑/LPIPS↓ TCLC-GS PSNR↑/SSIM↑/LPIPS↓
Scene-0008 25.32/0.82/0.26 26.41/0.85/0.22
Scene-0051 25.84/0.85/0.25 26.78/0.86/0.23
Scene-0058 24.71/0.81/0.27 26.45/0.88/0.20
Scene-0062 25.91/0.87/0.22 27.47/0.89/0.17
Scene-0129 26.87/0.88/0.21 28.65/0.90/0.16
Scene-0382 23.96/0.82/0.32 25.76/0.84/0.26
Average 25.44/0.84/0.26 26.92/0.87/0.21
Evaluation on Waymo Open Dataset: We evaluated the proposed method by comparing it with the baseline on the Waymo Open dataset. The performance comparison of image and depth synthesis from novel views, relative to the main baseline, is detailed in Table 1 and Table 2 separately. Additional comparisons of image synthesis from novel views against a broader range of baselines are depicted in Table 6. As indicated in Table 1, our method outperforms 3D-GS across individual scenes in terms of PSNR, SSIM, and LPIPS metrics for novel image synthesis. Furthermore, Table 6 shows that our method exceeds a broader array of baselines in average performance across these same metrics (PSNR, SSIM, and LPIPS) for novel image synthesis. Similarly, Table 2 demonstrates that our approach significantly surpasses 3D-GS in the depth synthesis metrics of AbsRel, RSME, and RSMElog. The significant improvement in depth synthesis performance can be attributed to the robust supervision provided by the rendered dense depths derived from the generated accurate 3D mesh. For a visual perspective, the comparison results of image synthesis from novel surrounding views are illustrated in Fig. 4. We can see that TCLC-GS renders more clear and accurate RGB images than the 3D-GS, especially on roadside objects and in distant areas viewed from the front, front-left, and front-right perspectives. Similarly, the visual comparison results for depth synthesis from novel surrounding views are showcased in Fig. 4. Here, TCLC-GS is observed to render denser and sharper depths compared to 3D-GS, especially in areas further away in the front, front-left, and front-right views.
Scene-0008 0.31/9.97/0.47 0.07/4.27/0.12
Scene-0051 0.47/8.68/0.60 0.05/2.21/0.07
Scene-0058 0.35/10.48/0.55 0.06/4.23/0.11
Scene-0062 0.32/8.22/0.39 0.04/4.03/0.08
Scene-0129 0.29/8.01/0.44 0.05/2.83/0.09
Scene-0382 0.32/10.91/0.51 0.05/4.58/0.10
Average 0.34/9.38/0.49 0.05/3.69/0.10
Evaluation on nuScenes Dataset: We conducted further comprehensive evaluations comparing our method with the baseline on the nuScenes dataset. The results of image and depth synthesis from novel viewpoints, compared with the primary baseline 3D-GS, are detailed in Table 3 and Table 4 separately. According to Table 3, our method outperforms 3D-GS in the metrics of PSNR, SSIM, and LPIPS for novel image synthesis. Likewise, as indicated in Table 4, our approach significantly excels beyond 3D-GS in the depth synthesis metrics of AbsRel, RSME, and RSMElog. The lower depth performance of our method observed on the nuScenes dataset, compared to the Waymo dataset, is due to the nuScenes dataset’s reliance on relatively sparse 32-line LiDAR data. In contrast, the Waymo dataset includes much denser LiDAR data, which contributes to a more accurate generation of 3D mesh and, subsequently, better rendered dense depth. In Table 6, we present a performance comparison between TCLC-GS and additional baselines on the nuScenes dataset, showcasing the average performance on PSNR, SSIM, and LPIPS metrics. Our method surpasses these baselines in all mentioned metrics, demonstrating its superior performance in image synthesis. Visual comparison results depicted in Figure 5 and 6 illustrate that our method not only produces clearer synthesized RGB images with fewer artifacts and less blurring but also generates sharper and denser synthesized depth images compared to the 3D-GS baseline.
@@
Method PSNR↑ SSIM↑ LPIPS↓
Mip-NeRF [1] 18.10 0.63 0.46
Mip-NeRF 360 [2] 23.45 0.74 0.35
S-NeRF [12] 25.62 0.77 0.27
3D-GS [11] 25.44 0.84 0.26
TCLC-GS 26.92 0.87 0.21